OpenStudy (theedwardsfamily):

Help is NEEDED fast!!!!! Solve. Express the answer in simplest form. 1/4+ w= 6 1/3 w=

OpenStudy (displayerror):

Convert the right hand side to an improper fraction: \[6 \frac{1}{3} = \frac{6 \times 3 + 1}{3} = \frac{19}{3}\] We can then rewrite the equation as \[\frac{1}{4} + w = \frac{19}{3}\] Multiply both sides by 3 and 4 in order to remove the fractions \[(3 \times 4) \times \left( \frac{1}{4} + w \right) = \frac{19}{3} \times (3 \times 4)\] This simplifies to \[3 + 12w = 76\] Which you should be able to solve on your own for w.

OpenStudy (theedwardsfamily):

@mathslover

OpenStudy (theedwardsfamily):

@DisplayError 7/4?

OpenStudy (displayerror):

Did you solve for w?

OpenStudy (theedwardsfamily):

yes. use calculator for some of my solving

OpenStudy (theedwardsfamily):

@DisplayError

OpenStudy (displayerror):

From the last equation that I typed up: \[3 + 12w = 76\] You used a calculator and got w = 7/4? Maybe you should solve it by hand instead.

OpenStudy (theedwardsfamily):

oh gtg! can u give me the right quick? @DisplayError

OpenStudy (theedwardsfamily):

the answer

OpenStudy (displayerror):

\[3 + 12w = 76\] \[12w = 76 -3\] \[12w = 73\] \[\frac{12w}{12} = \frac{73}{12}\]

OpenStudy (theedwardsfamily):

thanks

OpenStudy (theedwardsfamily):

is that the answer? @DisplayError

OpenStudy (theedwardsfamily):

help @sleepyjess !!!

OpenStudy (sleepyjess):

Do you understand what they are saying?

OpenStudy (theedwardsfamily):

nooo! need this done by MID-night! :(

OpenStudy (sleepyjess):

Ok, well first we need to make 1/4 and 1/3 have the same denominator

OpenStudy (sleepyjess):

Can you do that?

OpenStudy (theedwardsfamily):

yes

OpenStudy (sleepyjess):

Ok, so what do you get after you do that?

OpenStudy (theedwardsfamily):

find the LCD right?

OpenStudy (sleepyjess):

Yep

OpenStudy (sleepyjess):

What is the LCD?

OpenStudy (theedwardsfamily):

thats 1 right?

OpenStudy (theedwardsfamily):

@sleepyjess

OpenStudy (sleepyjess):

how would it be 1? what is 3*4?